# BRepFormer: Transformer-Based B-rep Geometric Feature Recognition

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; ccs: Computing methodologies Computer vision problemsccs: Applied computing Computer-aided manufacturingccs: Computing methodologies Shape modeling
@@ 1. Introduction
We propose BRepFormer, a transformer-based geometric feature recognition network that leverages a transformer architecture to effectively capture and process B-rep features. Unlike previous approaches that suffer from information loss and limited topological awareness (Ma et al., 2019), BRepFormer directly operates on the boundary representation (B-rep), ensuring high-fidelity feature extraction. Specifically, our model extracts both the geometric and topological features of the CAD model from multiple perspectives. During the feature encoding stage, geometric edge features and topological features are processed by separate encoders and then fused to form an attention bias, which serves as a constraint input into the transformer module. Meanwhile, the extracted geometric face features are encoded into tokens together with a virtual face. These tokens serve as carriers of information that are fed into the transformer module, facilitating deep interaction and information fusion among the features. Finally, the recognition head accepts the output from the transformer module and fuses the global and local features within it, achieving high-precision recognition of geometric features. By integrating these components, BRepFormer effectively propagates information across the CAD model structure, improving geometric feature recognition accuracy while preserving critical geometric and topological relationships.
We conducted experiments on the public MFInstSeg (Wu et al., 2024) and MFTRCAD (Xia et al., 2024) datasets, as well as our CBF dataset. The results demonstrate that BRepFormer achieves state-of-the-art recognition accuracy on these three datasets. Notably, BRepFormer excels in recognizing complex geometric structures and diverse machining features. On the MFTRCAD (Xia et al., 2024) dataset, it achieved a 93.16% recognition accuracy, surpassing the previous best method by 3.28 percentage points. To further validate the effectiveness of our approach, we performed detailed ablation studies, analyzing the impact of each model component on overall performance. The ablation study results indicate that each feature component contributes to an improvement in the model’s accuracy.
@@ 2.1. Rule-Based Geometric Feature Recognition
Rule-based geometric feature recognition methods identify geometric features in CAD models by applying predefined rules based on topological and geometric information. Early studies by Henderson (Henderson, 1984), Donaldson (Donaldson and Corney, 1993), and Chan (Chan and Case, 1994) introduced rule-based approaches that define boundary patterns and use expert systems for feature recognition. While intuitive and easy to implement, these methods struggle with flexibility due to the inherent limitations of predefined rules. Encoding all machining knowledge into a rule-based system is challenging, making maintenance cumbersome and leading to poor adaptability, particularly in handling complex intersecting features.
Volumetric decomposition is another well-established rule-based approach for geometric feature recognition. This method decomposes the material to be removed from a CAD model into intermediate volumes and reconstructs geometric features based on predefined rules. Woo et al. (Woo, 1982) introduced the Alternating Sum of Volumes (ASV) decomposition method, representing solids as a hierarchical structure of convex bodies through union and difference operations. Requicha et al. (Vandenbrande and Requicha, 1993) further developed an automatic feature recognition approach that decomposes machinable volumes into manufacturable features, addressing feature interactions using generation-testing strategies and computational geometry techniques. Wilde et al. (Kim and Wilde, 1992) refined the ASV method by introducing partitioning techniques, resolving convergence issues, and enhancing its applicability in feature recognition.
Graph-based methods have gained attention recently due to their strong alignment with the boundary representation (B-rep) structure of CAD models. These methods construct an Attribute Adjacency Graph (AAG) to capture the topological relationships and geometric attributes of faces and edges. Geometric features can then be identified by applying graph-matching techniques to detect subgraph patterns within the AAG. Joshi et al. (Joshi and Chang, 1988) first applied the AAG for topological and geometric information matching. Shah et al. (Gao and Shah, 1998) enhanced this approach by integrating hint-based feature recognition, introducing minimal condition subgraphs to improve the handling of feature interactions.
Hint-based methods (Han and Requicha, 1997; Li et al., 2015) offer a rule-driven approach for identifying complex intersecting features. They extract topological and geometric patterns, along with heuristic ”hints” derived from residuals left by feature intersections. These hints guide the reconstruction of incomplete feature information through reasoning. While this method improves the recognition of intersecting features and aids subsequent process planning, defining comprehensive and precise hints, along with robust reasoning rules, remains challenging. This complexity makes the hint-based method challenging to achieve fully automated and high-precision feature recognition in practical applications.
@@ 2.2. Deep Learning-Based Geometric Feature Recognition
Voxel-based approaches use 3D convolutional neural networks (CNNs) to process CAD models but face resolution-related information loss. FeatureNet (Zhang et al., 2018) applies 3D CNNs for feature recognition, while Peddireddy et al. (Peddireddy et al., 2020, 2021) refined voxelization techniques to predict machining processes like milling and turning. Despite these improvements, voxelization inherently reduces geometric fidelity, particularly at low resolutions. Mesh-based approaches seek to retain geometric details for improved recognition. Jia et al. (Jia et al., 2023) proposed an innovative method that combines the original MeshCNN with Faster RCNN, forming a geometric feature recognition scheme based on Mesh Faster RCNN. This approach enhances the accuracy of geometric feature detection while preserving the mesh geometry. However, the high memory demand when processing high-resolution data limits its application in large-scale scenarios.
In the CAD industry, B-rep (boundary representation) is the dominant format for 3D models, making graph-based representations particularly effective for geometric feature recognition. Graph neural networks (GNNs) have been widely applied due to their structural similarity with B-rep. Cao et al. (Cao et al., 2020) pioneered the transformation of B-rep models into graph-structured representations for learning. Colligan et al. (Colligan et al., 2022) introduced Hierarchical CADNet, a novel approach that leverages a two-level graph representation to improve recognition accuracy. Specifically, it utilizes the Face Adjacency Graph (FAG) to capture topological information and mesh patches to represent geometric details. Jayaraman et al. (Jayaraman et al., 2021) proposed UV-Net, which encodes surfaces and curves with CNNs and utilizes graph convolutional networks (GCNs) for feature learning.
@@ 3.2.1. Topological Feature Extraction
Face Angular Distance. To more accurately describe the relative position between two faces, our approach also extracts the dihedral angle between any two faces. When two faces share a common edge, the dihedral angle can explicitly indicate the geometric relationship formed by the two faces, whether it is concave, convex, or a smooth transition. For two non-adjacent faces, we use their normal vectors to calculate the angle between them. This method reflects the relative direction of the two faces in three-dimensional space, regardless of the series of intermediate faces through which they are connected. At the same time, our approach also constructs a matrix $M_{a}\in\mathbb{R}^{N\times N}$ to represent the angular information between faces, where each element $m_{a}(f_{i},f_{j})$ represents the angle between a face $f_{i}$ and another face $f_{j}$.
Face Centroid Distance. The centroid distance between faces, as another key topological feature, can be used to quantify their spatial separation by measuring the Euclidean distance between the centroids of two faces. To eliminate the influence of the model’s scale, we normalize this distance by the diagonal length of the bounding box of the entire solid model, obtaining a relative distance indicator. This normalized centroid distance can reflect the similarity between CAD models of different sizes and proportions. Similarly, we construct a matrix $M_{c}\in\mathbb{R}^{N\times N}$ to represent the angular information between faces, where each element $m_{c}(f_{i},f_{j})$ represents the Euclidean distance between the centroids of face $f_{i}$ and face $f_{j}$.
@@ 3.2.2. Geometric Feature Extraction
For the extraction of UV domain features from CAD models, our approach is inspired by UV-Net (Jayaraman et al., 2021). We sample and discretize the parametric surfaces and parametric curves surfaces in B-rep into regular 2D and 1D point grids with a uniform step size, as shown in Figure 2. For the grid feature representation of each face, the 3D coordinates, 3D normal vectors of the points, and an additional dimension indicating visibility are included. For the grid feature representation of each edge, the 3D coordinates and 3D tangent vectors of each point are also included. Moreover, to enhance the information representation, we add the normal vectors of the two adjacent faces of the edge to the edge grid points. Compared with discrete representation methods that may be insufficient in accurately describing complex surfaces and curves (such as voxel (Yan et al., 2023) grids and traditional meshes (Feng et al., 2019) ), our use of UV grids can capture precise geometric information and achieve a friendly input representation for neural networks.
@@ 4. Complex B-rep Feature (CBF) Dataset
We introduce a dataset named CBF to support research on complex geometric feature recognition, and describe the method for creating the corresponding dataset. The dataset comprises 20,000 CAD models in B-rep format. Each model is a B-rep model formed by boolean combinations of a base plate and three different geometric features on it. The faces of each geometric feature are labeled accordingly, with the label information stored in a separate JSON file to support research on model construction and training for these complex features.
@@ 5.1. Experimental Environment
We employed a single NVIDIA 4090 GPU and PyTorch-lightning V1.9.0 for network training, thereby emphasizing the lightweight characteristic of our network. In the training phase, AdamW was chosen as the optimizer, with an initial learning rate of 0.001. The parameters were set to $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon={1\times 10^{-8}}$ to guarantee training stability. Additionally, we employed the ReduceLROnPlateau (Al-Kababji et al., 2022) learning rate scheduling strategy, which dynamically adjusts the learning rate based on changes in validation loss. It is important to note that during the initial 5000 steps of training, we implemented a linear learning rate warm-up phase to help the model converge more effectively. The batch size was set to 64, and the entire training process lasted for a maximum of 200 epochs.
@@ 5.2. Experimental Datasets
We conducted experiments on the MFInstSeg and MFTRCAD public datasets to evaluate the machining feature recognition ability of our model and compared it with mainstream deep learning methods. The results indicate that our model achieved state-of-the-art accuracy on these datasets. Additionally, we verified the complex feature recognition capability of our model on the proposed CBF dataset. The results indicate that our network also achieves state-of-the-art accuracy. For all datasets, the data were split into 70% for training, 15% for validation, and 15% for testing.
@@ 5.3.1. Ablation Analysis of Geometric Features
The geometric features directly input into our BRepFormer  network include the UV domain geometric features and the attribute features of the B-rep. In this ablation study, a model with complete input features was used as the baseline, and then each of the three input features was removed one by one to generate the ablation models. The results shown in Table 6 indicate that the removal of any input feature leads to a decrease in feature recognition accuracy. Among them, the removal of attribute features causes the most significant performance drop, while the removal of UV domain geometric features has a less noticeable impact. This suggests that attribute features are more important than the input geometric features within our network architecture.

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 International Conference on Multimedia Retrieval; June 30-July 3, 2025; Chicago, IL, USAbooktitle: Proceedings of the 2025 International Conference on Multimedia Retrieval (ICMR ’25), June 30-July 3, 2025, Chicago, IL, USAdoi: 10.1145/3731715.3733283isbn: 979-8-4007-1877-9/2025/06ccs: Computing methodologies Computer vision problemsccs: Applied computing Computer-aided manufacturingccs: Computing methodologies Shape modeling
@@ 1. Introduction
We propose BRepFormer, a transformer-based geometric feature recognition network that leverages a transformer architecture to effectively capture and process B-rep features. Unlike previous approaches that suffer from information loss and limited topological awareness (Ma et al., 2019), BRepFormer directly operates on the B-rep, ensuring high-fidelity feature extraction. Specifically, our model extracts both the geometric and topological features of the CAD model from multiple perspectives. During the feature encoding stage, geometric edge features and topological features are processed by separate encoders and then fused to form an attention bias, which serves as a constraint input into the transformer module. Meanwhile, the extracted geometric face features are encoded into tokens together with a virtual face. These tokens serve as carriers of information that are fed into the transformer module, facilitating deep interaction and information fusion among the features. Finally, the recognition head accepts the output from the transformer module and fuses the global and local features within it, achieving high-precision recognition of geometric features. By integrating these components, BRepFormer effectively propagates information across the CAD model structure, improving geometric feature recognition accuracy while preserving critical geometric and topological relationships.
We conducted experiments on the public MFInstSeg (Wu et al., 2024) and MFTRCAD (Xia et al., 2024) datasets, as well as our CBF dataset. The results demonstrate that BRepFormer achieves state-of-the-art recognition accuracy on these three datasets. Notably, BRepFormer excels in recognizing complex geometric structures and diverse machining features. On the MFTRCAD (Xia et al., 2024) dataset, it achieved a 93.16% recognition accuracy, surpassing the previous best method by 3.28 percentage points. To further validate the effectiveness of our approach, we performed detailed ablation studies, analyzing the impact of each model component on overall performance. The results shows that every feature component helps boost the model’s accuracy.
@@ 2.1. Rule-Based Geometric Feature Recognition
Rule-based geometric feature recognition methods identify geometric features in CAD models by applying predefined rules based on topological and geometric information. Early studies by Henderson (Henderson, 1984), Donaldson (Donaldson and Corney, 1993), and Chan (Chan and Case, 1994) introduced rule-based approaches that define boundary patterns and use expert systems for feature recognition. However, these methods lack flexibility due to the limitations of predefined rules, making it difficult to encode all machining knowledge and adapt to complex features.
Volumetric decomposition is another well-established rule-based approach for geometric feature recognition. This method decomposes the material to be removed from a CAD model into intermediate volumes and reconstructs geometric features based on predefined rules. Woo et al. (Woo, 1982) introduced the Alternating Sum of Volumes (ASV) method, representing solids as hierarchical convex bodies. Requicha et al. (Vandenbrande and Requicha, 1993) further developed this approach to decompose machinable volumes into manufacturable features, addressing interactions through generation-testing strategies. Wilde et al. (Kim and Wilde, 1992) refined the ASV method by introducing partitioning techniques to enhance its applicability.
Graph-based methods have gained attention due to their alignment with the B-rep structure of CAD models. These methods construct an Attribute Adjacency Graph (AAG) to capture topological and geometric information, and identify features by detecting subgraph patterns. Joshi et al. (Joshi and Chang, 1988) first applied the AAG for topological and geometric information matching. Shah et al. (Gao and Shah, 1998) enhanced this approach by integrating hint-based feature recognition, introducing minimal condition subgraphs to improve the handling of feature interactions.
Hint-based methods (Han and Requicha, 1997; Li et al., 2015) use rules to identify complex intersecting features by extracting patterns and heuristic ”hints” from residuals left by feature intersections. These hints guide the reconstruction of incomplete feature information through reasoning. While this method improves the recognition of intersecting features and aids subsequent process planning, defining comprehensive and precise hints, along with robust reasoning rules, remains challenging. This complexity makes the hint-based method challenging to achieve fully automated and high-precision feature recognition in practical applications.
@@ 2.2. Deep Learning-Based Geometric Feature Recognition
Voxel-based approaches use 3D CNNs to process CAD models but suffer from resolution-related information loss. For example, FeatureNet (Zhang et al., 2018) applies 3D CNNs for feature recognition, and Peddireddy et al. (Peddireddy et al., 2020, 2021) refined voxelization techniques to predict machining processes like milling and turning. However, voxelization inherently reduces geometric fidelity, especially at low resolutions. Mesh-based methods aim to retain geometric details for better recognition. Jia et al. (Jia et al., 2023) proposed a method combining MeshCNN with Faster RCNN (Mesh Faster RCNN) to enhance detection accuracy while preserving mesh geometry. However, the high memory demand for processing high-resolution data limits its scalability in large-scale scenarios.
In the CAD industry, B-rep is the dominant format for 3D models, making graph-based representations particularly effective. Graph neural networks (GNNs) have been widely applied due to their structural similarity with B-rep. For example, Cao et al. (Cao et al., 2020) pioneered the transformation of B-rep models into graph-structured representations for learning. Colligan et al. (Colligan et al., 2022) introduced Hierarchical CADNet, a novel approach that leverages a two-level graph representation to improve recognition accuracy. Specifically, it utilizes the Face Adjacency Graph (FAG) to capture topological information and mesh patches to represent geometric details. Jayaraman et al. (Jayaraman et al., 2021) proposed UV-Net, which encodes surfaces and curves with CNNs and utilizes graph convolutional networks (GCNs) for feature learning.
@@ 3.2.1. Topological Feature Extraction
Face Angular Distance. To accurately describe the relative position between two faces, our approach extracts the dihedral angle between them. For adjacent faces sharing a common edge, the dihedral angle indicates their geometric relationship (e.g., concave, convex, or smooth transition). For non-adjacent faces, we calculate the angle between their normal vectors, reflecting their relative orientation in 3D space regardless of intermediate faces. At the same time, our approach also constructs a matrix $M_{a}\in\mathbb{R}^{N\times N}$ to represent the angular information between faces, where each element $m_{a}(f_{i},f_{j})$ represents the angle between a face $f_{i}$ and another face $f_{j}$.
Face Centroid Distance. The centroid distance between faces, a key topological feature, quantifies their spatial separation by measuring the Euclidean distance between face centroids. To eliminate scale influence, we normalize this distance by the bounding box diagonal length of the solid model, yielding a relative distance indicator. This normalized distance reflects similarity between CAD models of varying sizes and proportions. Similarly, we construct a matrix $M_{c}\in\mathbb{R}^{N\times N}$ to represent the angular information between faces, where each element $m_{c}(f_{i},f_{j})$ represents the Euclidean distance between the centroids of face $f_{i}$ and face $f_{j}$.
@@ 3.2.2. Geometric Feature Extraction
For extracting UV domain features from CAD models, we follow a method inspired by UV-Net (Jayaraman et al., 2021). We sample and discretize the parametric surfaces and parametric curves surfaces in B-rep into regular 2D and 1D point grids with a uniform step size, as shown in Figure 2. Each face’s grid includes 3D coordinates, normal vectors of the points, and a visibility indicator. Each edge’s grid includes 3D coordinates, tangent vectors, and the normal vectors of the adjacent faces. Compared to discrete methods like voxel grids (Yan et al., 2023) and traditional meshes (Feng et al., 2019), our UV grids capture precise geometric details and provide a suitable input format for neural networks.
@@ 4. Complex B-rep Feature (CBF) Dataset
We introduce the CBF dataset to support research on complex geometric feature recognition. It contains 20,000 B-rep CAD models, each formed by combining a base plate with three geometric features. The faces of these features are labeled and stored in JSON files for model construction and training research.
@@ 5.1. Experimental Environment
We trained our network using a single NVIDIA 4090 GPU and PyTorch-Lightning v1.9.0, highlighting its lightweight nature. During training, we used the AdamW optimizer with an initial learning rate of 0.001, and parameters set to $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon={1\times 10^{-8}}$ for stability. We also employed the ReduceLROnPlateau (Al-Kababji et al., 2022) learning rate scheduler to dynamically adjust the learning rate based on validation loss. A linear learning rate warm-up was implemented for the first 5000 steps to aid convergence. The batch size was set to 64, and training lasted up to 200 epochs.
@@ 5.2. Experimental Datasets
We evaluated our model’s machining feature recognition ability on the MFInstSeg and MFTRCAD public datasets and compared it with mainstream deep learning methods. Our model achieved state-of-the-art accuracy on these datasets. We also tested its complex feature recognition capability on the proposed CBF dataset, where it again achieved state-of-the-art accuracy. For all datasets, we used a 70% / 15% / 15% split for training, validation, and testing.
@@ 5.3.1. Ablation Analysis of Geometric Features
In our ablation study, we tested the importance of different input features for the BRepFormer network. The network takes UV domain geometric features and B-rep attribute features as inputs. We compared a baseline model with complete inputs to models that removed each of these features individually. The results in Table 6 show that removing any input feature reduces feature recognition accuracy. Specifically, removing attribute features causes the most significant performance drop, while removing UV domain geometric features has a smaller impact. This indicates that attribute features are more critical than geometric features in our network architecture.
@@ 6. Conclusion
7. Acknowledgments
This research was funded by the Fundamental Research Funds for the Central Universities, the Natural Science Basic Research Program of Shaanxi Province under Grant 2024JC-YBQN-0702, and NUS Research Scholarship and NUS ORIA.